# Sparse Model Soups: A Recipe for Improved Pruning via Model Averaging

### v1

@@ Abstract
Neural networks can be significantly compressed by pruning, leading to sparse models requiring considerably less storage and floating-point operations while maintaining predictive performance. Model soups (Wortsman et al., 2022a) improve generalization and out-of-distribution performance by averaging the parameters of multiple models into a single one without increased inference time. However, identifying models in the same loss basin to leverage both sparsity and parameter averaging is challenging, as averaging arbitrary sparse models reduces the overall sparsity due to differing sparse connectivities. In this work, we address these challenges by demonstrating that exploring a single retraining phase of Iterative Magnitude Pruning (IMP) with varying hyperparameter configurations, such as batch ordering or weight decay, produces models that are suitable for averaging and share the same sparse connectivity by design. Averaging these models significantly enhances generalization performance compared to their individual components. Building on this idea, we introduce Sparse Model Soups (SMS), a novel method for merging sparse models by initiating each prune-retrain cycle with the averaged model of the previous phase. SMS maintains sparsity, exploits sparse network benefits being modular and fully parallelizable, and substantially improves IMP’s performance. Additionally, we demonstrate that SMS can be adapted to enhance the performance of state-of-the-art pruning during training approaches.
@@ 1 Introduction
State-of-the-art Neural Network architectures typically rely on extensive over-parameterization with millions or billions of parameters (Zhang et al., 2016). In consequence, these models have significant memory requirements and the training and inference process is computationally demanding. However, recent work (LeCun et al., 1989; Hassibi and Stork, 1993; Han et al., 2015; Gale et al., 2019; Lin et al., 2020; Hoefler et al., 2021; Zimmer et al., 2022, 2023) has demonstrated that these resource demands can be significantly reduced by pruning the model, i.e., removing redundant structures such as individual parameters or groups thereof. The resulting sparse models demand considerably less storage and floating-point operations (FLOPs) during inference, while retaining performance comparable to dense models.
@@
A different line of research has shown that the performance of a predictor can be significantly enhanced by leveraging multiple models, instead of selecting the best one on a hold-out validation dataset and discarding the rest. Such ensembles combine the predictions of $m\in\mathbb{N}$ individually trained models by averaging their output predictions (Ganaie et al., 2021; Mehrtash et al., 2020; Chandak et al., 2023; Fort et al., 2019). Prediction ensembles have been shown to improve the predictive performance and positively impact predictive uncertainty metrics such as calibration, out-of-distribution generalization as well as model fairness (see e.g. Lakshminarayanan et al., 2017; Mehrtash et al., 2020; Allen-Zhu and Li, 2023; Ko et al., 2023). A significant drawback of ensembling is that all models have to be evaluated during deployment: the inference costs are hence increasing by a factor of $m$, a problem that has partially been addressed by leveraging an ensemble of sparsified, more efficient models (Liu et al., 2021; Whitaker and Whitley, 2022; Kobayashi et al., 2022).
@@
Figure 1: Constructing the average (middle) of two networks with different sparsity pattern (left, right) can reduce the overall sparsity level, turning pruned weights (dashed) into non-zero ones (solid). Weights reactivated in the averaged model are highlighted in orange.
@@
Several works propose to instead average the parameters (Izmailov et al., 2018; Wortsman et al., 2022a; Rame et al., 2022; Matena and Raffel, 2022), constructing a single model for inference. As opposed to prediction ensembles, which require sufficiently diverse models to enhance performance, such Model Soups (Wortsman et al., 2022a) require models that lie in a linearly connected basin of the loss landscape. Training two models from scratch with different random seeds will result in models whose parameter average will perform much worse than the individual models (Neyshabur et al., 2020) with recent research exploring neuron permutation to align them within a single basin (Singh and Jaggi, 2020; Ainsworth et al., 2023). Additionally, the computational advantages of sparse networks cannot be readily utilized, as averaging models with different sparse connectivities reduces the overall sparsity level (cf. Figure 1) and may require to prune again (Yin et al., 2022a, b), potentially resulting in further performance degradation.
@@
In this work, we tackle the challenge of concurrently leveraging sparsity as well as the benefits of combining multiple models into a single one. We draw our inspiration from recent work in the domain of transfer learning (Neyshabur et al., 2020; Wortsman et al., 2022a; Rame et al., 2022), which has shown that fine-tuning multiple copies of a pretrained model, differing only in random seed, yields models sufficiently similar for averaging and sufficiently diverse for performance improvements. At the core of our work lies the observation that a single prune-retrain phase in standard prune after training strategies, such as Iterative Magnitude Pruning (IMP) (Han et al., 2015), closely resembles the transfer learning paradigm. Starting from a pretrained model, the optimization objective shifts abruptly, either due to a new target domain or constraints imposed by the pruning subspace, followed by a subsequent training process termed ‘fine-tuning’, a term frequently used synonymously with ‘retraining’ to recover from pruning (Hoefler et al., 2021).
@@
We find that, akin to the fine-tuning phase in transfer learning, exploring the retraining phase after pruning with various hyperparameter configurations generates models that are amenable to averaging while sharing the same sparse connectivity by design. Such sparse averages exhibit superior performance compared to both their individual counterparts as well as to models retrained $m$ times as long. Taking it a step further, we initiate the subsequent prune-retrain cycle from the averaged model of the previous phase, which remarkably also enhances the performance of the individual retraining runs before averaging again. Our proposed approach, dubbed Sparse Model Soups (SMS), tackles the aforementioned challenges by enabling inference complexity independent of $m$, utilizes pretrained models without requiring training from scratch, preserves the sparsity pattern while leveraging sparse network benefits, and considerably improves IMP’s performance.
@@
    We demonstrate that pruning a well-trained model and retraining multiple copies with varied hyperparameter like batch ordering, weight decay, or retraining duration and length, produces models suitable for constructing an averaged model which exhibits superior generalization performance compared to its individual components. Importantly, these models maintain the same sparse connectivity of the pruned parent, preserved in their parameter average.
@@
    We propose Sparse Model Soups (SMS), a novel method for merging sparse models into a single classifier, leveraging the idea of starting each prune-retrain phase of IMP from an averaged model. SMS significantly enhances the performance of IMP in two ways: first, the average improves upon the individual models in terms of generalization performance and, secondly, the models retrained from an average exhibit better performance compared to those retrained from a single model.
@@
    We extend our findings to the pruning during training domain by incorporating SMS into magnitude-pruning based methods, yielding substantial performance improvements and enhancing their competitiveness in comparison with other state-of-the-art methods that sparsify during training.
@@
We start by formally presenting the context and our concept in Section 2. In Section 3, we back up our findings by providing extensive experimental results on image classification, semantic segmentation and neural machine translation architectures and datasets. We review related literature in Section 4 before concluding with a discussion in Section 5.
@@ 2 Methodology: Sparse Model Soups
In general, we represent neural network parameters as an $n$-dimensional vector $\theta\in\mathbb{R}^{n}$ and use the standard notation $[m]={\{1,\ldots,m\}}$ given $m\in\mathbb{N}$.
@@ 2.1 Preliminaries
Our focus lies on model pruning which aims at removing individual weights as exemplified by the previously introduced IMP approach. IMP, a prune after training algorithm, follows a three-stage pipeline. It starts with a pretrained model parameterized by $\theta$, prunes weights with magnitudes below a certain threshold, and then restores predictive power through retraining. This prune-retrain cycle is repeated multiple times, with each pruning step’s threshold determined by the suitable percentile to achieve the desired target sparsity after a predefined number of such phases. Recent studies (Gale et al., 2019; Zimmer et al., 2023) have demonstrated that magnitude pruning results in sparse models with performance competitive to much more complex algorithms.
Given $m$ sparse models $f_{\theta_{i}}$ with weights $\theta_{i}\in\mathbb{R}^{n},i\in[m]$, prediction ensembles construct a model as the functional equivalent of the average of the models’ output (see e.g. Liu et al., 2021; Whitaker and Whitley, 2022; Kobayashi et al., 2022). This ensemble requires $m$ forward passes for evaluation, but maintains the overall sparsity level. In contrast, our focus lies on examining the performance and sparsity of a single model, namely a linear combination of other models. Given scalars $\lambda_{i}\in\mathbb{R},i\in[m]$, we consider the prediction function $f_{\bar{\theta}}$, parameterized by the weights given by
$\bar{\theta}=\sum_{i\in[m]}\lambda_{i}\theta_{i}.$ (1)
@@ 2.2 Sparse Model Soups
Inspired by recent advancements in the transfer learning domain (Neyshabur et al., 2020; Wortsman et al., 2022a), which demonstrate that models fine-tuned from the same pretrained model albeit with different hyperparameter configurations end up in the same loss basin and can be combined into a soup, we hypothesize that a similar behavior can be achieved during fine-tuning or retraining from the same pruned model. Our motivation stems from the resemblance between the transfer learning paradigm and a single phase of IMP. When transitioning from the source to the target domain, the optimization objective changes abruptly, requiring adaptation (i.e., fine-tuning) to minimize the new objective. Similarly, ‘hard’ pruning alters the loss abruptly and requires adapting (i.e., retraining) given the newly added sparsity constraints.
@@
A single phase of this idea is illustrated on the left of Figure 2. The pretrained model $\theta$ is pruned, yielding model $\theta_{p}$, which is then replicated $m$ times. Subsequently, each of the $m$ models is independently retrained with different hyperparameter configurations, such as varying random seeds, weight decay factors, retraining lengths, or learning rate schedules. Finally, the $m$ retrained models are merged into a single model. This process ensures that all $m$ retrained models $\theta_{1},\ldots,\theta_{m}$ share the same sparsity pattern, as they all originate from the same pruned network. However, when combining models after multiple prune-retrain cycles, identical sparsity connectivity between all models is not guaranteed. To address this, we average the models after each phase and commence the subsequent phase with the previously averaged model. The resulting method, termed Sparse Model Soups (SMS), is presented as pseudocode on the right of Figure 2.
@@
SMS offers several benefits and addresses key challenges. First, the inference complexity of the final model remains independent of $m$. The method is highly modular, allowing for different hyperparameter configurations and different $m$ in each phase. Further, the retraining of the $m$ models can be fully parallelized, enhancing efficiency. By initiating each phase with the merged model from the previous one, sparsity patterns are preserved, and the advantages of sparse networks are utilized; as the number of cycles increases, the networks become sparser, potentially leading to further efficiency gains. Moreover, SMS effectively leverages the benefits of large pretrained models without the need for training from scratch.
Merging strategies
In practice, effectively combining models with improved generalization can be challenging. For instance, when using a wide range of weight decay values, models may end up far apart, and merging them could result in decreased performance. Although Equation 1 permits any linear combination of models, we primarily experiment with two convex combination methods proposed by Wortsman et al. (2022a): UniformSoup and GreedySoup. The former assigns an equal factor $\lambda_{i}=1/m$ to each model $i\in[m]$. In contrast, GreedySoup sorts models by their validation accuracy, starting with the best-performing model and progressively adding models from the sorted list only if their inclusion enhances the validation accuracy compared to the previous subset.
@@ 3 Experimental results
We first outline our general experimental approach, including datasets, architectures, and metrics. To ensure comparability and standardization in the rapidly evolving field of pruning, we follow the guidelines proposed by Blalock et al. (2020) whenever possible. Our implementation is based on the PyTorch framework (Paszke et al., 2019), using Weights & Biases (Biewald, 2020) for logging. To enable reproducibility, our code is available at github.com/ZIB-IOL/SMS. We evaluate our approach on well-known datasets for image recognition, semantic segmentation, and neural machine translation (NMT), including ImageNet-1K (Russakovsky et al., 2015), CIFAR-10/100 (Krizhevsky et al., 2009), Celeb-A (Liu et al., 2015), CityScapes (Cordts et al., 2016), and WMT16 DE-EN (Bojar et al., 2016). We utilize state-of-the-art architectures, such as ResNets (He et al., 2015), WideResNets (Zagoruyko and Komodakis, 2016), MaxViT (Tu et al., 2022), PSPNet (Zhao et al., 2017), and the T5 transformer (Raffel et al., 2020). We primarily analyze the tradeoff between model sparsity and test performance, using accuracy, mean Intersection over Union (mIoU), and BLEU score as main evaluation metrics. The appendix contains detailed hyperparameters, standard training setups and metric definitions. We also consider the sparsity-induced theoretical speedup (Blalock et al., 2020) and fairness assessed by per-group-recalls as secondary measures. For validation, we use 10% of the training data. Weight magnitude serves as the criterion for unstructured pruning, while for convolutional filters, we follow Li et al. (2016) and prune using filter norms. For retraining, we stick to the linear learning rate schedules proposed by Zimmer et al. (2023).
Recomputing Batch-Normalization (BN) (Ioffe and Szegedy, 2015) statistics is crucial in both pruning and model averaging, as observed by Li et al. (2020) and Jordan et al. (2022), respectively. When reporting test or validation accuracy for single or averaged models, we reset all BN layers and recompute statistics using a forward pass on the entire training dataset. Notably, we fixed the batch order of the dataloader during recomputation, as even random batching could skew results, with more details in the appendix.
@@ 3.1 Evaluating Sparse Model Soups
Table 1: WideResNet-20 on CIFAR-100 and ResNet-50 on ImageNet: Test accuracy comparison of SMS to several baselines for target sparsities 98% (top) and 90% (bottom) given three prune-retrain cycles. We report results using UniformSoup as well as GreedySoup for merging. Results are averaged over multiple seeds with standard deviation included. The best value is highlighted in bold.
@@
We evaluate SMS against several key baselines, comparing it at each prune-retrain phase to the best-performing single model among all candidates for averaging (best candidate), the mean candidate, regular IMP (i.e., $m=1$) and an extended version of IMP retrained $m$ times as long (IMP+). Unlike SMS, IMP+ is not parallelizable in each phase, with runtime directly dependent on $m$. Moreover, we contrast SMS with executing regular IMP $m$ times, then averaging the models post-final phase (termed IMP-AVG), addressing potential sparsity reduction by re-pruning the model (Yin et al., 2022b). For SMS, we vary random seeds for each of the $m$ models, with additional results in the appendix. In all phases except the first, candidates for averaging, including the best one, are retrained from a previous model soup, distinguishing best candidate from regular IMP, which does not rely on averaging.
@@
Table 1 presents results for three-phase IMP using WideResNet-20 on CIFAR-100 and ResNet-50 on ImageNet. The left column lists the baselines, while the three main columns correspond to phases and sparsities, targeting 98% sparsity for CIFAR-100 and 90% for ImageNet. Each main column has three subcolumns, indicating the number of models to average (3, 5, or 10). For full results on different architectures, datasets, target sparsities, and structured pruning, we refer to the appendix. We find that SMS consistently improves upon the test accuracy of the best candidate, often with a 1% or higher margin. This confirms that the models after retraining are averageable, resulting in better generalization than individual models. SMS notably improves upon both classical IMP and its extended retraining variant, IMP+, with up to 2% enhancements even when using $m=3$ splits.
However, remarkably, we find that the best candidates in phases two and three are also consistently better than IMP and often than IMP+, even though they are not part of the average yet. While some improvement is expected when selecting the best model from multiple candidates, the fact that the average test accuracy of all candidates (i.e., mean candidate) also surpasses IMP and IMP+ suggests that starting from a model soup, rather than just a retrained model from the previous phase as classical IMP does, enhances generalization throughout the subsequent phase. This is surprising since the better-performing model soup is more affected by pruning, which further supports the generalization-stability tradeoff (Bartoldson et al., 2020): pruning’s generalization benefit increases with more instability to pruning, as measured by pruning-induced performance degradation. IMP-AVG naively averages models generated by IMP in the final phase. This often leads to reduced sparsity because of differing sparse connectivities (cf. Figure 1), requiring a subsequent pruning that tends to degrade performance compared to individual models.
@@
In summary, we find that averaging after each phase and starting subsequent phases from the soup of the previous phase capitalizes on two dynamics, which often enable significant improvements over IMP: first of all, the model soup consistently improves upon individual soup candidates, demonstrating that pruned and retrained models are indeed averageable, resulting in better generalization performance. Secondly, models retrained from a pruned soup also outperform those following the classical prune-retrain cycle. Pruning a model with higher generalization performance yields better models after retraining, despite experiencing a larger pruning-induced performance drop.
@@
Table 1 also distinguishes between the uniform and greedy selection strategies. When varying only the random seed for each of the $m$ models, none appear to converge to a different basin, making it unnecessary to include only a subset in the soup. We observe that the uniform approach usually outperforms the greedy one in terms of both SMS and the best candidate metric: in the last phase of CIFAR-100, the best and mean uniform candidate substantially outperform the best greedy candidate, indicating that previous greedy soups yield less performant models after retraining.

### v3

@@ Abstract
Neural networks can be significantly compressed by pruning, yielding sparse models with reduced storage and computational demands while preserving predictive performance. Model soups (Wortsman et al., 2022a) enhance generalization and out-of-distribution (OOD) performance by averaging the parameters of multiple models into a single one, without increasing inference time. However, achieving both sparsity and parameter averaging is challenging as averaging arbitrary sparse models reduces the overall sparsity due to differing sparse connectivities. This work addresses these challenges by demonstrating that exploring a single retraining phase of Iterative Magnitude Pruning (IMP) with varied hyperparameter configurations such as batch ordering or weight decay yields models suitable for averaging, sharing identical sparse connectivity by design. Averaging these models significantly enhances generalization and OOD performance over their individual counterparts. Building on this, we introduce Sparse Model Soups (SMS), a novel method for merging sparse models by initiating each prune-retrain cycle with the averaged model from the previous phase. SMS preserves sparsity, exploits sparse network benefits, is modular and fully parallelizable, and substantially improves IMP’s performance. We further demonstrate that SMS can be adapted to enhance state-of-the-art pruning-during-training approaches.
@@ 1 Introduction
State-of-the-art Neural Network architectures typically rely on extensive over-parameterization with millions or billions of parameters (Zhang et al., 2016). In consequence, these models have significant memory requirements and the training and inference process is computationally demanding. However, recent work (e.g. Han et al., 2015; Lin et al., 2020; Renda et al., 2020; Zimmer et al., 2022) has demonstrated that these resource demands can be significantly reduced by pruning the model, i.e., removing redundant structures such as individual parameters or groups thereof. The resulting sparse models demand considerably less storage and floating-point operations (FLOPs) during inference, while retaining performance comparable to dense models.
@@
A different line of research has shown that the performance of a predictor can be significantly enhanced by leveraging multiple models, instead of selecting the best one on a hold-out validation dataset and discarding the rest. Such ensembles combine the predictions of $m\in\mathbb{N}$ individually trained models by averaging their output predictions (Ganaie et al., 2021; Mehrtash et al., 2020; Chandak et al., 2023; Fort et al., 2019). Prediction ensembles have been shown to improve the predictive performance and positively impact predictive uncertainty metrics such as calibration, out-of-distribution generalization as well as model fairness (Lakshminarayanan et al., 2017; Mehrtash et al., 2020; Allen-Zhu & Li, 2023; Ko et al., 2023). A significant drawback of ensembling is that all models have to be evaluated during deployment: the inference costs are hence increasing by a factor of $m$, a problem that has partially been addressed by leveraging an ensemble of sparsified, more efficient models (Liu et al., 2021; Whitaker & Whitley, 2022; Kobayashi et al., 2022).
@@
Figure 1: Creating the average (middle) of two networks with different sparsity patterns (left, right) may lower overall sparsity, changing pruned weights (dashed) to non-zero (solid), with reactivated weights highlighted in orange.
@@
Several works propose to instead average the parameters (Izmailov et al., 2018; Wortsman et al., 2022a; Rame et al., 2022; Matena & Raffel, 2022), constructing a single model for inference. Unlike prediction ensembles that require sufficiently diverse models for better performance, such Model Soups (Wortsman et al., 2022a) need models to lie in a linearly connected basin of the loss landscape. However, training models from scratch with differing random seeds but identical initialization often yields models whose parameter average will perform much worse than the individual models (Neyshabur et al., 2020) with recent studies investigating neuron permutation to align them within a single basin (Singh & Jaggi, 2020; Ainsworth et al., 2023). Beyond the initial challenge of identifying networks suitable for averaging, another problem emerges when attempting to leverage the computational advantages of sparse networks: averaging models with different sparse connectivities reduces overall sparsity (cf. Figure 1) and may require to prune again (Yin et al., 2022a; b), potentially resulting in further performance degradation.
@@
In this work, we tackle the challenge of concurrently leveraging sparsity as well as the benefits of combining multiple models into a single one. We draw our inspiration from recent work in the domain of transfer learning (Neyshabur et al., 2020; Wortsman et al., 2022a; Rame et al., 2022), which has shown that fine-tuning multiple copies of a pretrained model, differing only in random seed, yields models sufficiently similar for averaging and sufficiently diverse for generalization improvements. At the core of our work lies the observation that a single prune-retrain phase in standard prune after training strategies, such as Iterative Magnitude Pruning (IMP, Han et al., 2015), closely resembles the transfer learning paradigm. Starting from a pretrained model, the optimization objective shifts abruptly, either due to a new target domain or subspace constraints imposed by pruning, followed by a training process termed ‘fine-tuning’, often used interchangeably with ‘retraining’ to recover from pruning (Hoefler et al., 2021).
@@
We find that, akin to the fine-tuning phase in transfer learning, exploring various hyperparameter configurations during the retraining phase after pruning generates models that are suitable for averaging while sharing the same sparse connectivity by design. Such sparse averages exhibit superior performance compared to both their individual counterparts as well as to models retrained $m$ times as long, effectively reducing IMP’s runtime. Additionally, we initiate the next prune-retrain cycle from the averaged model just obtained, which remarkably also enhances the performance of the individual retraining runs before averaging again. Our proposed approach, Sparse Model Soups (SMS), tackles the aforementioned challenges and enables inference complexity independent of $m$, utilizes pretrained models without requiring training from scratch, preserves the sparsity pattern while leveraging sparsity benefits, and considerably improves IMP’s generalization and OOD performance.
@@
    We demonstrate that pruning a well-trained model and retraining multiple copies with varied hyperparameter like batch ordering, weight decay, or retraining duration and length, produces models suitable for constructing an averaged model which exhibits superior generalization and OOD performance compared to its individual components. Importantly, these models retain the sparsity pattern of their pruned parent, preserved in their parameter average.
@@
    We propose Sparse Model Soups (SMS), a novel method for merging sparse models into a single classifier, leveraging the idea of starting each prune-retrain phase of IMP from an averaged model. SMS significantly enhances the performance of IMP in two ways: first, the average improves upon the individual models in terms of generalization and OOD performance and, secondly, the models retrained from an average exhibit better performance compared to those retrained from a single model.
@@
    We extend our findings to the pruning during training domain, demonstrating SMS’s versatility by integrating it with multiple other state-of-the-art approaches. This yields substantial performance improvements and enhances their competitiveness in comparison to other leading methods that sparsify during training.
@@
We introduce our framework in Section 2. In Section 3, we experimentally validate our findings across image classification, semantic segmentation, and neural machine translation architectures and datasets. Section 4 reviews relevant literature, followed by a discussion in Section 5.
@@ 2.1 Preliminaries
Our focus lies on model pruning which aims at removing individual weights as exemplified by the previously introduced IMP approach. IMP, a prune after training algorithm, follows a three-stage pipeline. It starts with a pretrained model parameterized by $\theta$, prunes weights with magnitudes below a certain threshold, and then restores predictive power through retraining. This prune-retrain cycle is repeated multiple times, with each pruning step’s threshold determined by the suitable percentile to achieve the desired target sparsity after a predefined number of such phases. Recent studies (Gale et al., 2019; Zimmer et al., 2023) have demonstrated that magnitude pruning results in sparse models with performance competitive to more complex algorithms.
Given $m$ sparse models $f_{\theta_{i}}$ with weights $\theta_{i}\in\mathbb{R}^{n},i\in{\{1,\ldots,m\}}$, prediction ensembles construct a model as the functional equivalent of the average of the models’ output (Liu et al., 2021; Whitaker & Whitley, 2022; Kobayashi et al., 2022). This ensemble requires $m$ forward passes for evaluation, but maintains the overall sparsity level. In contrast, our focus lies on examining the performance and sparsity of a single model, specifically a linear combination of other models. Given scalars $\lambda_{i}\in\mathbb{R}$, we consider the prediction function $f_{\bar{\theta}}$, parameterized by the weights given by
$\bar{\theta}=\sum_{1\leq i\leq m}\lambda_{i}\theta_{i}.$ (1)
@@ 2.2 Sparse Model Soups
Inspired by recent advancements in the transfer learning domain (Neyshabur et al., 2020; Wortsman et al., 2022a), which demonstrate that models fine-tuned from the same pretrained model end up in the same loss basin and can be combined into a soup, we hypothesize that a similar behavior can be achieved during retraining from the same pruned model. Our motivation stems from the resemblance between the transfer learning paradigm and a single phase of IMP. When transitioning from the source to the target domain, the optimization objective changes abruptly, requiring adaptation (i.e., fine-tuning) to minimize the new objective. Similarly, ‘hard’ pruning alters the loss abruptly and requires adapting (i.e., retraining) given the newly added sparsity constraints.
@@
A single phase of this idea is illustrated on the left of Figure 2. The pretrained model’s weights $\theta$ are pruned, yielding model $\theta_{p}$, which is then replicated $m$ times. In this setup, pruned weights remain permanently non-trainable. Subsequently, each of the $m$ models is independently retrained with different hyperparameter configurations, such as varying random seeds, weight decay factors, retraining lengths, or learning rate schedules. Finally, the $m$ retrained models are merged into a single model. This process ensures that all $m$ retrained models $\theta_{1},\ldots,\theta_{m}$ share the same sparsity pattern, as they all originate from the same pruned network with a fixed pruning mask. However, when combining models after multiple prune-retrain cycles, identical sparsity connectivity between all models is not guaranteed. To address this, we average the models after each phase and begin the subsequent phase with the previously averaged model. The resulting method, termed Sparse Model Soups (SMS), is presented as pseudocode on the right of Figure 2.
@@
SMS offers several benefits and addresses key challenges. First, the inference complexity of the final model remains independent of $m$. The method is highly modular, allowing for different hyperparameter configurations and different $m$ in each phase. Further, the retraining of the $m$ models can be fully parallelized, enhancing efficiency as detailed in Section 3.2. By initiating each phase with the merged model from the previous one, sparsity patterns are preserved, and the advantages of sparse networks are utilized; as the number of cycles increases, the networks become sparser, potentially leading to further efficiency gains. Moreover, SMS effectively leverages the benefits of large pretrained models without the need for training from scratch.
Effectively merging models for enhanced generalization can be challenging, as models may end up far apart. We primarily employ two convex combination methods from Wortsman et al. (2022a): UniformSoup and GreedySoup. UniformSoup equally weighs each model with $\lambda_{i}=1/m$. On the other hand, GreedySoup orders models by validation accuracy, sequentially including models only if they improve validation accuracy over the prior subset.
@@ 3 Experimental results
We first outline our general experimental approach. For reproducibility, our implementation is available at github.com/ZIB-IOL/SMS. We evaluate our approach on well-known datasets for image recognition, semantic segmentation, and neural machine translation (NMT), including ImageNet-1K (Russakovsky et al., 2015), CIFAR-10/100 (Krizhevsky et al., 2009), Celeb-A (Liu et al., 2015), CityScapes (Cordts et al., 2016), WMT16 DE-EN (Bojar et al., 2016) and the benchmark OOD-datasets CIFAR-100-C and ImageNet-C (Hendrycks & Dietterich, 2019). We utilize state-of-the-art architectures, such as ResNets (He et al., 2015), WideResNets (Zagoruyko & Komodakis, 2016), MaxViT (Tu et al., 2022), PSPNet (Zhao et al., 2017), and the T5 transformer (Raffel et al., 2020). For validation, we use 10% of the training data. We use magnitude-based unstructured pruning and filter norm-based structured pruning as suggested by Li et al. (2016). For retraining, we stick to the linear learning rate schedules LLR and ALLR (Zimmer et al., 2023), with further details in Section A.2. Appendix A describes exact hyperparameters and settings for pretraining, pruning and retraining.
Recomputing Batch-Normalization (BN) (Ioffe & Szegedy, 2015) statistics is crucial in both pruning and model averaging, as observed by Li et al. (2020) and Jordan et al. (2022), respectively. When reporting test accuracy for single or averaged models, we reset all BN layers and recompute statistics using a forward pass on the entire training dataset.
@@ 3.1 Evaluating Sparse Model Soups
Table 1: WideResNet-20 on CIFAR-100 and ResNet-50 on ImageNet: Test accuracy comparison for target sparsities 98% (top) and 90% (bottom) given three prune-retrain cycles. We report results using UniformSoup as well as GreedySoup for merging. Results are averaged over multiple seeds with standard deviation included. The best value is highlighted in bold.
@@
[…]
Table 1 presents results for three-phase IMP using WideResNet-20 on CIFAR-100 and ResNet-50 on ImageNet, employing ALLR and ten retraining epochs per phase. For SMS, we vary the random seeds across each of the $m$ models. The three main columns correspond to phases and sparsities, targeting 98% sparsity for CIFAR-100 and 90% for ImageNet. Each main column has three subcolumns, indicating the number of models to average (3, 5, or 10). We discuss the main observations below.
  1. 1.
@@
    SMS significantly enhances generalization. We find that SMS consistently improves upon the test accuracy of the best candidate, often with a 1% or higher margin. This confirms that the models after retraining are averageable, resulting in better generalization than individual models. SMS notably improves upon both regular IMP and its extended retraining variant, $\text{IMP}_{m\!\times}$, with up to 2% enhancements even when using $m=3$ splits.
  2. 2.
    Starting from a model soup enhances generalization. Surprisingly, the best candidates in the second and third phase frequently exceed both IMP and $\text{IMP}_{m\!\times}$. While some improvement is anticipated when picking the best among multiple candidates, it is notable that the mean candidate accuracy (i.e., mean candidate) often surpasses both IMP and $\text{IMP}_{m\!\times}$ as well. This suggests that initiating from a soup, as opposed to starting from a singular model as in regular IMP, enhances generalization in the subsequent phase.
  3. 3.
    IMP-RePrune faces sparsity reduction and performance degradation. Naively averaging IMP’s models in the final phase often leads to reduced sparsity due to differing sparse connectivities, requiring repruning which typically degrades performance compared to individual models. Under certain conditions, IMP-RePrune can remain competitive, suggesting that similar pruning patterns may emerge across multiple pruning rounds (cf. Section B.1).
@@
In summary, we find that averaging after each phase and starting subsequent phases from the soup of the previous phase capitalizes on two dynamics, which often enable significant improvements over IMP: first of all, the model soup consistently improves upon individual soup candidates, demonstrating that pruned and retrained models are indeed averageable and exhibit enhanced generalization. Secondly, models retrained from a pruned soup also outperform those following the classical prune-retrain cycle. Pruning a model with higher generalization performance yields better models after retraining, despite experiencing a larger pruning-induced performance drop. For full results on different architectures, datasets, target sparsities, and structured pruning, we refer to Section B.1.
@@
Table 1 also contrasts uniform and greedy soup selections. With just the random seed varied for the $m$ models, none appear to diverge to a different basin, rendering greedy subset selection unnecessary. The uniform approach predominantly outperforms the greedy one, notably when comparing the best or mean candidates in scenarios like the last phase of CIFAR-100, indicating that retraining from previous greedy soups yields less performant models.